WHAT is 60 + 60 x 0
Answer by addingup(3677)

Follow the rules of operations. First you multiply, then add.
If you wanted to add and then multiply you'd have to write it like this: (60+60)*0 because the rule is that you clear the parenthesis first.
Since you do not have a parenthesis:
60*0=0+60= 60 This is your answer.